OpenStudy (anonymous):

A parachutist of mass 75kg descends with an acceleration of 8 m/s2. What is the upward force exerted by the air on the parachutist? How would you go about finding this answer?

OpenStudy (anonymous):

Forces acting are :- a) Weight, Wg b) upward air resistance, Fr F.net = Wg - Fr => 75*9.8 - Fr = 75*8 => Fr = 75*(9.8 - 8) => Fr = 135 N

OpenStudy (anonymous):

drfigure out the forces ====find the net thats all just like shubham has done